# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v4

@@ 3. Classification of orbits
Lemma 3.6. Let $\lambda\in S,D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1). Suppose that for every $\alpha\in D$, the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$, and formula (2) holds for all $\gamma^{\prime}\succ\gamma$ that satisfy condition (1). Then formula (2) also holds for $\lambda$, $D$ and $\gamma$.
Proof. Consider
$(\exp(\chi).\lambda)(e_{\gamma})=\sum\limits_{\alpha\in D}\lambda(e_{\alpha})T_{\gamma,\alpha}(x)=\lambda(e_{\gamma})+\sum\limits_{\alpha\in D,\alpha\succ\gamma}\lambda(e_{\alpha})T_{\gamma,\alpha}(x).$
@@
Let $\alpha\in D$ with $\alpha\succ\gamma$. Since $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. By the definition of $S$, the root $\alpha$ satisfies condition (1). Therefore, by the inductive hypothesis, $\lambda(e\alpha)$ can also be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
$\lambda(e_{\alpha})=(\exp(\chi).\lambda)(e_{\alpha})-F_{D,\alpha}((\exp(\chi).\lambda)(e_{\beta}),\beta\succ\alpha).$
@@
Consequently, the difference $(\exp(\chi).\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ is a rational function of $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, which means condition (2) is satisfied. $\square$
We are now ready to prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$.
@@
Proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. If $D=\{\alpha\}$, then Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ holds by Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Indeed, in this case $(\exp(\chi).\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$. Moreover, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$. Finally, for $\gamma\prec\alpha$ satisfying condition (1), the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
@@
Now we will prove formula (2) step by step for all $\lambda\in S$ by considering the supports from the list one after another. (Note that most supports can be treated similarly to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$; see below.) Recall that $\lambda=\sum\limits_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum\limits_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi).\lambda\in M^{*}$.
@@
  • Suppose the largest root (with respect to the lexicographic order) in the support is smaller than $\alpha_{2}+\alpha_{3}+\alpha_{4}$. For all such supports and for every $\gamma\in\Phi^{+}$ satisfying condition (1), Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, so the proof is complete. (Indeed, for all relevant $\gamma$, the polynomials $T_{\gamma,\alpha}(x)$ are, up to a non-zero scalar factor, equal to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0, or 11.)
@@
  • Suppose the largest root in the support equals $\alpha_{2}+\alpha_{3}+\alpha_{4}$. There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$. Since Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ covers the case of one-element supports, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$). The conditions of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ are also satisfied, so nothing remains to prove in this case.
@@
  • Next, suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}$. Here the first nontrivial case appears, i.e., Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be proved for all supports of this type by a straightforward application of Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Such a support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$. The complication arises because the table entries at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ both involve the same root $\alpha_{2}+\alpha_{3}$. However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$. Consequently, a situation may occur where $\alpha_{4}$ satisfies condition (1) but Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ cannot be applied directly. Therefore, in this case we must compute the coordinates of $\mu$ explicitly:
    $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}=&~\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=&~\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}}=&-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-\\ &-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{3}+\alpha_{4}}=&-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&-x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{3}}=&-\lambda([x,e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{3}}]])=-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{2}}e_{\alpha_{2}},[x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{4}}e_{\alpha_{4}},e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x_{\alpha_{4}}e_{\alpha_{4}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}}])\\ =&-C_{\alpha_{2}+\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\dfrac{1}{2}C_{\alpha_{2},\alpha_{3}}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\\ &+\dfrac{1}{2}(C_{\alpha_{4},\alpha_{3}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}+C_{\alpha_{2},\alpha_{3}}C_{\alpha_{4},\alpha_{2}+\alpha_{3}})x_{\alpha_{2}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&~2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}+x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}-x_{\alpha_{2}}x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{4}}=&~\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{4}}])+\frac{1}{2}\lambda([x,[x,e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})+\\ &+\dfrac{1}{2}C_{\alpha_{3},\alpha_{4}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\\ &+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\mu_{\alpha_{3}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
@@
  • Now suppose the largest root in the support equals either $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$. In these cases, the statement also follows almost trivially from Lemmas Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ and Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. The only nontrivial situation occurs for the support $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$. At first glance, it might seem that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$. However, this is not the case. From Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}$. Consequently, $x_{\alpha_{3}}$ can be recovered from $(\exp(\chi).\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$. Since $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ depends only on $x_{\alpha_{3}}$.
  • If the largest root in the support equals $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the nontrivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$. This case follows immediately from the formulas in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ together with the analysis of the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
@@
  • Next, if the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those containing the roots $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3}+\alpha_{4}$. One can easily show that
@@
    For the second support, the equations for the coordinates of $\mu$ are:"
@@
  • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$. Although there are many such supports, they are all amenable to a simple analysis. Each such support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, and $\alpha_{1}+\alpha_{2}+\alpha_{3}$. Consequently, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$, and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as rational function of $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$). Moreover, $x_{\alpha_{1}}$ can be expressed as a rational function of $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$. It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ appears in a monomial in a coordinate $\mu_{\alpha}$, then that monomial can be expressed as a rational function of the coordinates $\mu_{\beta}$ with $\beta\succ\alpha$. Furthermore, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$. Finally, those monomials that cannot be expressed in this way satisfy formula (2) by the analysis above. A similar analysis applies to supports whose largest root is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
The remaining steps can be analyzed similarly: either Lemma Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ applies, or the equations for the coordinates can be written out explicitly in a straightforward manner. Thus, Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ is proved. $\square$

### v5

@@ 3 Classification of orbits
Let $\lambda\in S,D=\operatorname{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1). Suppose that for every $\alpha\in D$, the pair $(\gamma,\alpha)$ belongs to $\zeta^{\prime}_{D,\gamma}$, and formula (2) holds for all $\gamma^{\prime}\succ\gamma$ that satisfy condition (1). Then formula (2) also holds for $\lambda$, $D$ and $\gamma$.
Proof 3.8.
Consider
$(\exp(\chi)\cdot\lambda)(e_{\gamma})=\sum_{\alpha\in D}\lambda(e_{\alpha})T_{\gamma,\alpha}(x)=\lambda(e_{\gamma})+\sum_{\alpha\in D,\alpha\succ\gamma}\lambda(e_{\alpha})T_{\gamma,\alpha}(x).$
@@
Let $\alpha\in D$ with $\alpha\succ\gamma$. Since $(\gamma,\alpha)\in\zeta^{\prime}_{D,\gamma}$, the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$. By the definition of $S$, the root $\alpha$ satisfies condition (1). Therefore, by the inductive hypothesis, $\lambda(e\alpha)$ can also be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, because
$\lambda(e_{\alpha})=(\exp(\chi)\cdot\lambda)(e_{\alpha})-F_{D,\alpha}((\exp(\chi)\cdot\lambda)(e_{\beta}),\beta\succ\alpha).$
@@
Hence, the difference $(\exp(\chi)\cdot\lambda)(e_{\gamma})-\lambda(e_{\gamma})$ is a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\gamma}$, which means condition (2) is satisfied.
We are now ready to prove Theorem 3.2.
@@
Proof of Theorem 3.2. If $D=\{\alpha\}$, then Theorem 3.2 holds by Lemma 3.4. Indeed, in this case $(\exp(\chi)\cdot\lambda)(e_{\gamma})=\lambda(e_{\alpha})\cdot T_{\gamma,\alpha}(x)$. Moreover, $T_{\gamma,\alpha}(x)=0$ for $\gamma\succ\alpha$, while $T_{\gamma,\alpha}(x)=1$ for $\gamma=\alpha$. Finally, for $\gamma\prec\alpha$ satisfying condition (1), the polynomial $T_{\gamma,\alpha}(x)$ can be expressed as a rational function of $\{(\exp(\chi)\cdot\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
@@
Now we will prove formula (2) step by step for all $\lambda\in S$ by considering the supports from the list one after another. (Note that most supports can be treated similarly to Example 3.3; see below.) Recall that
$\lambda=\sum_{\beta\in\Phi^{+}}\lambda_{\beta}e_{\beta}^{*}\in\mathfrak{n}^{*},~\mu=\sum_{\beta\in\Phi^{+}}\mu_{\beta}\otimes e_{\beta}^{*}=\exp(\chi)\cdot\lambda\in M^{*}$.
@@
  • Suppose the largest root (with respect to the lexicographic order) in the support is smaller than $\alpha_{2}+\alpha_{3}+\alpha_{4}$. For all such supports and for every $\gamma\in\Phi^{+}$ satisfying condition (1), Lemma 3.7 applies, so the proof is complete. (Indeed, for all relevant $\gamma$, the polynomials $T_{\gamma,\alpha}(x)$ are, up to a non-zero scalar factor, equal to $x_{\alpha_{2}}$, $x_{\alpha_{3}}$, $x_{\alpha_{4}}$, 0, or 11.)
@@
  • Suppose the largest root in the support equals $\alpha_{2}+\alpha_{3}+\alpha_{4}$. There are only two such supports in the list: $\{\alpha_{2}+\alpha_{3}+\alpha_{4}\}$ and $\{\alpha_{2}+\alpha_{3}+\alpha_{4},\alpha_{3}\}$. Since Lemma 3.4 covers the case of one-element supports, the pair $(\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4})$ belongs to $\zeta^{\prime}_{D,\alpha_{3}}$ regardless of whether $\alpha_{3}$ lies in $D$ (the presence of simple roots in the support does not affect $\zeta^{\prime}_{D,\gamma}$). The conditions of Lemma 3.7 are also satisfied, so nothing remains to prove in this case.
@@
  • Next, suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}$. Here the first nontrivial case appears, i.e., Theorem 3.2 cannot be proved for all supports of this type by a straightforward application of Lemma 3.7. Such a support may also contain $\alpha_{2}+\alpha_{3}+\alpha_{4}$ and $\alpha_{4}$. The complication arises because the table entries at positions $(\alpha_{3},\alpha_{2}+2\alpha_{3})$ and $(\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4})$ both involve the same root $\alpha_{2}+\alpha_{3}$. However, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}}(x)\neq T_{\alpha_{4},\alpha_{2}+\alpha_{3}+\alpha_{4}}(x)$. Consequently, a situation may occur where $\alpha_{4}$ satisfies condition (1) but Lemma 3.7 cannot be applied directly. Therefore, in this case we must compute the coordinates of $\mu$ explicitly:
    $\begin{split}\mu_{\alpha_{2}+2\alpha_{3}}=&~\lambda(e_{\alpha_{2}+2\alpha_{3}})=\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=&~\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{2}+\alpha_{3}}=&-\lambda([x,e_{\alpha_{2}+\alpha_{3}}])=\\ =&-\lambda([x_{\alpha_{4}}e_{\alpha_{4}}+x_{\alpha_{3}}e_{\alpha_{3}},e_{\alpha_{2}+\alpha_{3}}])=-C_{\alpha_{4},\alpha_{2}+\alpha_{3}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})-\\ &-C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})=x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}-2x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}},\\ \mu_{\alpha_{3}+\alpha_{4}}=&-\lambda([x,e_{\alpha_{3}+\alpha_{4}}])=-\lambda([x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}+\alpha_{4}}])=-C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&-x_{\alpha_{2}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{3}}=&-\lambda([x,e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x,[x,e_{\alpha_{3}}]])=-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{3}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{2}}e_{\alpha_{2}},[x_{\alpha_{3}}e_{\alpha_{3}}+x_{\alpha_{4}}e_{\alpha_{4}},e_{\alpha_{3}}])+\dfrac{1}{2}\lambda([x_{\alpha_{4}}e_{\alpha_{4}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{3}}])\\ =&-C_{\alpha_{2}+\alpha_{3}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\dfrac{1}{2}C_{\alpha_{2},\alpha_{3}}C_{\alpha_{3},\alpha_{2}+\alpha_{3}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+2\alpha_{3}})+\\ &+\dfrac{1}{2}(C_{\alpha_{4},\alpha_{3}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}+C_{\alpha_{2},\alpha_{3}}C_{\alpha_{4},\alpha_{2}+\alpha_{3}})x_{\alpha_{2}}x_{\alpha_{4}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})\\ =&~2x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}+x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+2\alpha_{3}}-x_{\alpha_{2}}x_{\alpha_{4}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}},\\ \mu_{\alpha_{4}}=&~\lambda(e_{\alpha_{4}})-\lambda([x,e_{\alpha_{4}}])+\frac{1}{2}\lambda([x,[x,e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-\lambda([x_{\alpha_{2}+\alpha_{3}}e_{\alpha_{2}+\alpha_{3}},e_{\alpha_{4}}])+\\ &+\dfrac{1}{2}\lambda([x_{\alpha_{3}}e_{\alpha_{3}},[x_{\alpha_{2}}e_{\alpha_{2}},e_{\alpha_{4}}]])=\lambda_{\alpha_{4}}-C_{\alpha_{2}+\alpha_{3},\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})+\\ &+\dfrac{1}{2}C_{\alpha_{3},\alpha_{4}}C_{\alpha_{2},\alpha_{3}+\alpha_{4}}x_{\alpha_{2}}x_{\alpha_{3}}\lambda(e_{\alpha_{2}+\alpha_{3}+\alpha_{4}})=\lambda_{\alpha_{4}}+x_{\alpha_{2}+\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\\ &+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}}\lambda_{\alpha_{2}+\alpha_{3}+\alpha_{4}}=\lambda_{\alpha_{4}}+\dfrac{\mu_{\alpha_{2}+\alpha_{3}}\mu_{\alpha_{3}+\alpha_{4}}-\mu_{\alpha_{2}+\alpha_{3}+\alpha_{4}}\mu_{\alpha_{3}}}{2\mu_{\alpha_{2}+2\alpha_{3}}}.\end{split}$
@@
  • Now suppose the largest root in the support equals $\alpha_{2}+2\alpha_{3}+\alpha_{4}$ or $\alpha_{2}+2\alpha_{3}+2\alpha_{4}$. In these cases, the statement also follows almost trivially from Lemma  3.4 and Lemma 3.7. The only nontrivial situation occurs for the support
    $D=\{\alpha_{2}+2\alpha_{3}+2\alpha_{4},\alpha_{2}+2\alpha_{3},\alpha_{2}\}$.
    At first glance, it might seem that $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ cannot be expressed as a rational function in $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}}$. However, this is not the case. From Lemma 3.4 we know that $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$ can be expressed as a rational function of
    $\{(\exp(\chi)\cdot\lambda)(e_{\beta})\}_{\beta\succ\alpha_{2}+\alpha_{3}+\alpha_{4}}.$
    Consequently, $x_{\alpha_{3}}$ can be recovered from $(\exp(\chi)\cdot\lambda)(e_{\alpha_{2}+\alpha_{3}})$ and $T_{\alpha_{2}+\alpha_{3},\alpha_{2}+2\alpha_{3}+2\alpha_{4}}(x)$. Since $T_{\alpha_{2},\alpha_{2}+2\alpha_{3}}(x)$ depends only on $x_{\alpha_{3}}$.
  • If the largest root in the support equals $\alpha_{1}$ or $\alpha_{1}+\alpha_{2}$, then the nontrivial case is when $D$ contains $\alpha_{1}+\alpha_{2},\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}$. This case follows immediately from the formulas in Example 2.7 together with the case $D=\{\alpha_{2}+2\alpha_{3},\alpha_{2}+\alpha_{3}+\alpha_{4}\}$.
@@
  • Next, if the largest root in the support equals $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, then the most interesting supports are those containing the roots
    $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4},\quad\alpha_{2}+2\alpha_{3}+2\alpha_{4}\quad\text{and}\quad\alpha_{2}+2\alpha_{3}+\alpha_{4}.$
    One can easily show that
@@
    For the second support, the equations for the coordinates of $\mu$ are:
@@
  • Suppose the largest root in the support equals $\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}$. Although there are many such supports, they are all amenable to a simple analysis. Each such support does not contain the roots $\alpha_{1}+\alpha_{2}+2\alpha_{3}$, $\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}$, and $\alpha_{1}+\alpha_{2}+\alpha_{3}$. Consequently, $x_{\alpha_{4}}$, $x_{\alpha_{3}}$, and $x_{\alpha_{3}+\alpha_{4}}$ can always be expressed as rational function of $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}+\alpha_{4}}$, $\mu_{\alpha_{1}+\alpha_{2}+\alpha_{3}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ (in fact, as polynomials with coefficients from $\overline{\mathbb{F}}[\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}^{-1}]$). Moreover, $x_{\alpha_{1}}$ can be expressed as a rational function of $\mu_{\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ and $\mu_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$. It follows that if $\lambda_{\alpha_{1}+\alpha_{2}+2\alpha_{3}+\alpha_{4}}$ appears in a monomial in a coordinate $\mu_{\alpha}$, then that monomial can be expressed as a rational function of the coordinates $\mu_{\beta}$ with $\beta\succ\alpha$. Furthermore, almost all other monomials can also be expressed as polynomials in $x_{\alpha_{4}},~x_{\alpha_{3}},~x_{\alpha_{3}+\alpha_{4}}$ and $x_{\alpha_{1}}$. Finally, those monomials that cannot be expressed in this way satisfy formula (2) by the analysis above. A similar analysis applies to the cases where the largest root of the support is $2\alpha_{1}+3\alpha_{2}+4\alpha_{3}+2\alpha_{4}$.
The remaining steps can be analyzed similarly: either Lemma 3.7 applies, or the equations for the coordinates can be written out explicitly in a straightforward manner. Thus, Theorem 3.2 is proved. $\square$